OpenStudy (anonymous):

The length of one leg of a right triangle is 24 ft, and the length of the hypotenuse is 40 ft. What is the length of the other leg?

OpenStudy (whpalmer4):

 Remember the Pythagorean theorem: \[a^2+b^2=c^2\]

OpenStudy (anonymous):

so would the answer be 6.3?

OpenStudy (whpalmer4):

If 6.3^2 + 24^2 = 40^2, that would be correct. Is it?